OpenStudy (anonymous):

write an equation in slope-inyercept form for the line with slope -4 that contains the points(0,-3)

OpenStudy (anonymous):

Point-slope form: (y+3)=-4(x) y=-4x-3

OpenStudy (anonymous):

slope intercept form of a line is y = m.x + c where m is the slope and c is the intercept of the line. now you know the slope (= -4). all you want to know is the intercept. Since point (0,-3) is in the line so the equation of line will satisfy for the given line i.e, -3 = -4 (0) + c => c = -3 now we have both slope and intercept, so the equation of line is y = -4x -3 give me medal, if i really helped you. :)